OpenStudy (anonymous):

Sally was asked to find two integers such that one is three less than 4 times the other and the sum of the two integers is 42. The integers must be _____. A. 14, 28 B. 15, 27 C. 9, 33 D. -13, 55

OpenStudy (anonymous):

It's C.

OpenStudy (anonymous):

9*4= 36 36-3= 33 9+33=42